OpenStudy (anonymous):

Which expressions show how 8 • 54 can be rewritten using the distributive property? 8 • 50 + 4 8 • 50 + 8 • 4 8 • 60 – 8 • 6 8 • 60 + 8 • 4

OpenStudy (anonymous):

B because \[8×(54)=8×(5+4)=8×(5)+8×(4)\]